1.)Solve using the multiplication principle: -3/4x=-5/6
2.) Find the slope, if it exists, of the line containing the pair of points:
(9,3) and (10,-1)

Answer by rfer(16322)

1)
-3/4x=-5/6
x=-5/6*-4/3
x=20/18
x=1 1/9
2)
slope m=-4